OpenStudy (anonymous):

A=B+3(4-C) If B equals 5 and C equals 2, what is the value of A?

OpenStudy (gw2011):

By substituting in the values you get: A=5+3(4-2) A=5+3(2) A=5+6 A=11

OpenStudy (anonymous):

11